# Beyond Calibration: Assessing the Probabilistic Fit of Neural Regressors via Conditional Congruence

### v2

@@ 3.2 Conditional Congruence Error
Unless otherwise specified, we use CLIP [38] as our choice of $\varphi$ due to its impressive representational capacity and ease of implementation using libraries such as OpenCLIP [39]. For $k_{\mathcal{X}}$, we use the polynomial kernel, $k(\mathbf{x},\mathbf{x}^{\prime})=(\frac{1}{d}\mathbf{x}^{T}\mathbf{x}^{\prime}+1)^{3}$, similar to [40]. When forming $\mathcal{S}^{\prime}_{\varphi}$, we draw $\ell=1$ samples per test point. For $k_{\mathcal{Y}}$, we use the radial basis function (RBF) kernel $k(\mathbf{x},\mathbf{x}^{\prime})=\exp{-\gamma\left\lVert\mathbf{x}-\mathbf{x}^{\prime}\right\rVert^{2}}$, with $\gamma=\frac{1}{2\hat{\sigma_{y}^{2}}}$, where $\hat{\sigma_{y}^{2}}=\frac{1}{n-1}\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}$ is estimated from our test targets. Empirically, we find that this choice of kernels and hyperparameters yields well-behaved CCE values that match with intuition. For a more extensive investigation of hyperparameters, see Appendix C.3.
Once we fix $\mathcal{S}$, the CCE is defined for any input $x\in\mathcal{X}$, even if we do not have an associated label $y\in\mathcal{Y}$. This makes it a powerful tool for quantifying model reliability in a variety of deployment settings where labels are not known. We show how this property enables point-wise diagnostics in Section 4.3, and demonstrate how it yields a natural algorithm for the subfield of machine learning with rejection in Appendix C.1.
To evaluate a model’s congruence, we compute the CCE for each $x_{i}$ in our test data. The resultant array of values provides a notion of probabilistic alignment across the entire input space. When a single summary statistic is desired, we take the mean of these CCE values (denoted $\overline{\mbox{CCE}}$). An algorithm describing our process in greater detail can be found in Appendix B.1.
@@
(a) Gaussian Blur perturbation to the marginal $P_{X}$.
(b) Label noise perturbations to the marginal $P_{Y}$.
(c) Mixup perturbations to the conditional $P_{Y|X}$.
Figure 3: We train three neural regression models (Poisson, $\beta$-Gaussian [41], and DDPN [16]) on the count regression task of COCO-People and apply progressively increasing levels of perturbation to the test data. We assess the ability of CCE to describe the misalignment of each model’s predictive distribution. (a) We increase the level of Gaussian blur ($\sigma$) to the pixels of each image, thus modifying $P_{X}$; (b) we apply Gaussian noise with constant variance ($\mu,\sigma=1.0$) to each label, modifying $P_{Y}$; (c) We apply the Mixup augmentation [42], $(1-\alpha)z_{1}+\alpha z_{2}$, to both the images and labels and perturb the conditional, $P_{Y|X}$. Under all three perturbation regimes, we see that CCE is effective at describing increasing levels of incongruence between the model’s predictive distribution and the test data. This effect is invariant to the choice of model.
@@ 4 EXPERIMENTS
To demonstrate the effectiveness of CCE, we show how it can both quantify a model’s overall probabilistic fit and act as a point-wise diagnostic tool for identifying specific model shortcomings. First, we show how CCE effectively recovers the true discrepancy between a predictive distribution and test data when the data generating process is known. Second, we show how CCE scales to high-dimensional image regression tasks, and again effectively identifies distributional misalignment under known perturbations to the data. Third, we show how CCE compares to existing measures of probabilistic fit across a variety of image regression tasks. Finally, we include two case studies in which we use the CCE to pinpoint areas in the input space where models are most- and least-congruent. All code used in these experiments is freely available online [^1] https://anonymous.4open.science/r/probcal-8294.
In our experiments, CCE is compared to ECE, which directly measures calibration, and Negative Log Likelihood (NLL), which is a proper scoring rule [43]. In contrast to ECE, NLL can be evaluated point-wise, but requires distributional assumptions, is difficult to compare across regressors, and is sensitive to outliers. In Appendix D.2 we show that point-wise NLL is subject to these pathologies and that CCE avoids them. Finally, in Appendix C.3 we perform extensive hyperparameter studies for $\lambda$, $\gamma$, $\ell$ and the choice of kernel.
4.1 CCE with a Known Data-Generating Process (Synthetic Data)
@@
Our primary aim is to demonstrate that CCE can effectively evaluate a model’s overall probabilistic fit. We first use a synthetic dataset [16]) where the true data-generating process is known. The data describe a one-dimensional discrete synthetic regression problem that exhibits various levels of dispersion, with high levels of uncertainty at low counts and severe under-dispersion at high counts. Both continuous and discrete neural regressors can be fit to this dataset. We train a small multi-layer perceptron (MLP) to output the parameters of a Poisson, Negative Binomial (NB), Gaussian, and Double Poisson distribution [16] using the appropriate NLL loss. The models’ predictive distributions over the test split of the dataset are visualized in row 1 of Figure 2, with the respective NLL and ECE of the model overlaid on the plot. We also compute the CCE for each model on all test inputs. We plot the resultant values in row 2 of Figure 2 (with $\overline{\mbox{CCE}}$ indicated in each respective plot). Since our inputs, $x_{i}$, are one-dimensional, we set $\varphi$ to be the identity and use the RBF kernel with $\gamma=0.5$ for $k_{\mathcal{X}}$.
@@
The Poisson and Negative Binomial distributions do not have the flexibility to represent under-dispersion. Accordingly, they struggle most in the $[\pi,2\pi]$ region of the input (where the requisite regression targets have very low variance). This is where the CCE identifies the largest incongruence between the model and the data. Meanwhile, the Gaussian and Double Poisson models achieve much more aligned predictive distributions. This is supported by their near-zero CCE values. Measuring congruence via CCE allows us to (rightly) conclude that the Gaussian NN and Double Poisson NN achieve the best probabilistic fit.
In Figure 2 we observe disagreement between the CCE and ECE when assessing the probabilistic fit of the Gaussian and Double Poisson models. In fact, as further expanded upon in Appendix D.1, this may be due to an implicit bias in the ECE computation against discrete models. We also note that though the NLL suggests a similar ranking of models as the CCE in this scenario, its undetermined lower bound makes it difficult to quantify how close to perfectly fit each model is, since the NLL cannot be easily interpreted as an estimate of $D_{C}(f,P_{Y|X})$.
Table 1: Comparison of ECE, NLL, and $\overline{\mbox{CCE}}$. We denote the best performer on each metric in bold and the second-best performer with an underline.
COCO-People FG-Net AAF
ECE ($\downarrow$) NLL ($\downarrow$) $\overline{\mbox{CCE}}$ ($\downarrow$) ECE ($\downarrow$) NLL ($\downarrow$) $\overline{\mbox{CCE}}$ ($\downarrow$) ECE ($\downarrow$) NLL ($\downarrow$) $\overline{\mbox{CCE}}$ ($\downarrow$)
Poisson NN 0.088 2.316 0.061 (0.001) 0.039 3.027 0.062 (0.011) 0.035 3.324 0.026 (0.003)
NB NN 0.067 2.248 0.072 (0.002) 0.023 2.863 0.043 (0.014) 0.023 3.380 0.022 (0.003)
DDPN 0.052 2.270 0.075 (0.001) 0.054 3.300 0.082 (0.010) 0.091 3.433 0.051 (0.004)
Gaussian NN 0.078 2.411 0.125 (0.001) 0.050 3.144 0.049 (0.001) 0.027 3.432 0.026 (0.003)
Stirn et al. [12] 0.081 2.650 0.140 (0.002) 0.031 3.142 0.065 (0.019) 0.104 4.230 0.042 (0.002)
Immer et al. [13] 0.057 2.298 0.081 (0.001) 0.626 10.843 0.657 (0.001) 0.026 3.430 0.031 (0.003)
Seitzer et al. [41] 0.055 2.374 0.072 (0.002) 0.093 3.618 0.079 (0.003) 0.115 4.252 0.063 (0.001)
Table 2: Comparison of ECE, NLL, and $\overline{\mbox{CCE}}$ for the EVA dataset (a continuous image regression task). We denote the best performer on each metric in bold and the second-best performer with an underline.

### v3

@@ 3.2 Conditional Congruence Error (CCE)
Unless otherwise specified, we use CLIP [38] as our choice of $\varphi$ due to its impressive representational capacity and ease of implementation using libraries such as OpenCLIP [20]. For $k_{\mathcal{X}}$, we use the polynomial kernel $k(\mathbf{x},\mathbf{x}^{\prime})=(\mathbf{x}^{T}\mathbf{x}^{\prime}+1)^{3}$, similar to Bińkowski et al. [4]. To ensure boundedness of the polynomial kernel, we apply L2-normalization to the outputs of $\varphi$ (thus constraining the inner product to live in [1,1][-1,1]). When forming $\mathcal{S}^{\prime}_{\varphi}$, we draw $\ell=1$ samples per test point. For $k_{\mathcal{Y}}$, we use the radial basis function (RBF) kernel $k(\mathbf{x},\mathbf{x}^{\prime})=\exp{-\gamma\left\lVert\mathbf{x}-\mathbf{x}^{\prime}\right\rVert^{2}}$, with $\gamma=\frac{1}{2\hat{\sigma_{y}^{2}}}$, where $\hat{\sigma_{y}^{2}}=\frac{1}{n-1}\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}$ is estimated from our test targets. Empirically, we find that this choice of kernels and hyperparameters yields well-behaved CCE values that match with intuition. For a more extensive investigation of hyperparameters, see see Section 5.
Once we fix $\mathcal{S}$, the CCE is defined for any input $x\in\mathcal{X}$, even if we do not have an associated label $y\in\mathcal{Y}$. This makes it a powerful tool for quantifying model reliability in a variety of deployment settings where labels are not known. We show how this property enables point-wise diagnostics in Section 4.7, and demonstrate how it yields a natural algorithm for the subfield of machine learning with rejection in Appendix A.3.
To evaluate a model’s congruence, we compute CCE for each $x_{i}$ in our test data. The resultant array of values provides a notion of probabilistic alignment across the entire input space. When a single summary statistic is desired, we take the mean of these CCE values (denoted $\overline{\mbox{CCE}}$). An algorithm describing our process in greater detail can be found in Appendix C.1.
@@ 4 Experiments
To demonstrate the effectiveness of CCE, we identify four properties that an ideal estimator of probabilistic alignment should have. We then evaluate CCE across these dimensions and compare to common measures such as ECE and NLL. The ideal properties we seek to confirm in CCE are: 1) Correctness - the estimator can accurately characterize the true discrepancy between the learned predictive distribution and the distribution that generated the data; 2) Monotonicity - as correctness worsens, the magnitude of the estimator increases; 3) Reliability - the estimator consistently produces similar values across multiple evaluations; 4) Robustness - the estimator is not unduly influenced by a small set of points.
First, in Section 4.3 we show how CCE satisfies the correctness property using a synthetic dataset where detected misalignment can be visually confirmed (and ECE and NLL exhibit limitations). Second, we establish that CCE displays monotonicity on complex image regression tasks by applying progressively increasing levels of perturbations to the test data (Section 4.4). Third, we demonstrate the reliability of CCE by performing a large-scale study of seven deep regression models across four datasets (Section 4.5). We isolate sources of variability and demonstrate that CCE exhibits no worse reliability than ECE. Fourth, we show that CCE is more robust to outliers than NLL (Section 4.6). Finally, we include a case study where we use CCE to pinpoint areas in the input space where predictions are most and least congruent, without ground truth labels at the evaluation points. All code used in these experiments is freely available online. [^1] https://github.com/spencermyoung513/probcal
4.1 Datasets
In our experiments, we perform several image regression tasks where the goal is to accurately predict either a count (discrete regression) or a real number (continuous regression). COCO-People is a dataset constructed from the ‘person’ class of the MS-COCO dataset [29] where the bounding boxes for each image have been translated into a single, discrete count [46]. Both FG-Net [36] and AAF [6] are image datasets where the task is defined as predicting a person’s age from an image of their face. Lastly, EVA [24] is an image dataset describing how appealing an image is on a continuous scale [0, 10].
4.2 Baselines
Estimators of Probablistic Fit
In our experiments, we compare CCE to ECE [26], a popular measure of calibration, and Negative Log Likelihood (NLL), a well-known proper scoring rule [15]. ECE is solely an aggregate statistic, while NLL can be computed point-wise.
Deep Regression Models
On all counting tasks (COCO-People, FG-Net, and AAF), we train a Poisson DNN [14], a Negative Binomial DNN [45], DDPN [46], a vanilla Gaussian NN [34], and modern Gaussian regressors such the $\beta$-modified Gaussian ($\beta=0.5$) [40], the naturally parameterized Gaussian [21], and the “faithful” heteroscedastic regressor [42]. For the EVA dataset, we only benchmark Gaussian models since the labels are continuous. Detailed specifications for these experiments can be found in Appendix B.2.
Figure 2: We test the correctness property of CCE by visualizing the probabilistic fit of four DNNs trained on a synthetic dataset with discrete regression targets. (Row 1) A plot of the predicted mean and 95% credible interval for each model, along with its ECE. (Row 2) NLL incurred on the test points, along with the overall mean value. (Row 3) CCE between the test points and predictive distributions. We plot the mean and one standard deviation range at each point, estimated with a bootstrap approximation of the sampling distribution of CCE. ECE and NLL standard errors are also computed with the bootstrap. CCE is able to accurately characterize the quality of each model’s predictive distribution, relative to the test data.
4.3 Evaluating the correctness of CCE
@@
We first investigate if CCE can correctly characterize a model’s overall probabilistic fit. To do so, we use a synthetic dataset where the true data-generating process is known [46]. The data describe a one-dimensional discrete regression problem that exhibits various levels of dispersion, with high levels of uncertainty at low counts and severe under-dispersion at high counts. We train a small multi-layer perceptron (MLP) to output the parameters of a Poisson, Negative Binomial (NB), Gaussian, and Double Poisson distribution [46] using the appropriate NLL loss. The models’ predictions over the test split of the dataset are visualized in row 1 of Figure 2, with the respective ECE of each model overlaid on the plot. Row 2 depicts point-wise NLL values for each model on the test set. In row 3, we plot CCE values, along with a bootstrapped approximation of the standard error. We set $\varphi=I$ and use the RBF kernel with $\gamma=0.5$ for $k_{\mathcal{X}}$.
@@
The Poisson and Negative Binomial distributions cannot represent under-dispersion. Accordingly, they struggle most in the $\pi,2\pi]$ region of the input, where the labels have very low variance. This is where CCE identifies the largest incongruence between the model and the data. Meanwhile, the Gaussian NN and DDPN models achieve better-aligned predictive distributions. This is supported by their near-zero CCE values, allowing us to (rightly) conclude that these models achieve the best fit.
While average NLL yields a similar ranking of models as CCE, its pointwise evaluations are less informative, with sensitivity to outliers and less identifiability of poorly-fit regions (see Section 4.6 for further discussion of this property). Finally, we note that CCE and ECE disagree when assessing the fit of the Gaussian NN and DDPN. As further expanded upon in Appendix D.1, this is likely due to to an implicit bias in the ECE computation against discrete models.
COCO-people
AAF
(a) Gaussian Blur ($\tilde{P}_{X}$)
(b) Label noise ($\tilde{P}_{Y}$)
(c) Mixup $\tilde{P}_{Y|X}$.
(d) Gaussian Blur ($\tilde{P}_{X}$)
(e) Label noise ($\tilde{P}_{Y}$)
(f) Mixup $\tilde{P}_{Y|X}$.
Figure 3: We test the monotonicity property of CCE. We measure its ability to describe the misalignment between the model’s predictive distribution and the test data under progressively larger perturbations. For both the COCO-People and AAF regression datasets, we train three DNNs. We analyze the behavior of CCE under Gaussian blur, label noise, and mixup corruptions. In general, CCE increases monotonically as the corruptions become more severe. We run 5 trials for each perturbation and report the mean and standard deviation (error bars indicate $\pm 1$ standard deviation).
4.4 Evaluating the monotonicity of CCE
We now demonstrate that CCE is monotonic with respect to underlying changes in the distribution of the input, $X$, and labels, $Y$. To do so, we perform a series of image regression tasks and show that CCE detects increasing divergence between a model’s predictive distribution and the test data under progressively larger perturbations.
For each model, we perform three types of data shift. First, we apply varying levels of Gaussian blur ($\sigma$) to the input image, which perturbs the marginal of the input, $P_{X}$, away from the training data. Second, we apply Gaussian noise with constant variance, $\epsilon\sim\mathcal{N}(\mu,1)$, to the marginal of the output, $P_{Y}$. Third, we apply the Mixup transform, $(1-\alpha)z_{1}+\alpha z_{2}$, to both the images and labels, which is a perturbation to the conditional, $P_{Y|X}$ [48]. We denote the perturbed distributions as $\tilde{P}_{X}$, $\tilde{P}_{Y}$ and $\tilde{P}_{Y|X}$, respectively. For each model, we perform five trials of the perturbation and evaluation, and plot the mean and standard deviation of the resulting $\overline{\mbox{CCE}}$ statistic. For each test point, we draw a single MC sample from the predictive distribution ($\ell=1$).
Results are shown in Figure 3. We see that under all three data shift regimes, $\overline{\mbox{CCE}}$ smoothly increases as perturbations of progressively larger magnitude are applied. Additionally, we observe that because CCE is a measure of conditional congruence, it can identify misalignment in both the conditional, $P_{Y|X}$, and marginal distributions, $P_{X}$ and $P_{Y}$. We therefore conclude that CCE satisfies monotonicity.
Table 1: We test the reliability property of CCE by performing a two-sample t-test (with both a two-sided and a one-sided alternative) comparing the mean coefficient of variation ($\mu_{CV}$) of CCE to ECE. Both tests yield an insignificant $p$-value, suggesting that the reliability of CCE is no different from that of popular existing metrics like ECE.
$H_{a}$ $p$-value
$\mu_{CV}^{(CCE)}>\mu_{CV}^{(ECE)}$ 0.811
$\mu_{CV}^{(CCE)}\neq\mu_{CV}^{(ECE)}$ 0.378
4.5 Evaluating the reliability of CCE
Figure 4: We test the robustness property of CCE, compared to NLL. From left to right: A t-SNE projection of the test split of AAF (color indicates age), the CCE values achieved by a trained Gaussian DNN, and the NLL values from this model on each test input. NLL is sensitive to outliers, highlighting a handful of points (see (-40, -18) and (-10, 10)) as being exceptionally poorly fit while assigning roughly equal values to all other points. Meanwhile, CCE illuminates a broader set of points with faulty predictions, shedding light on the local structure of congruence. In this case, many of regions of high incongruence appear to line up with images of older faces, suggesting the model cannot reliably assess the ages of elderly people.
In this section, we seek to probe the reliability of CCE. There are two potential sources of variability in CCE: 1) randomness due to the Monte Carlo (MC) sampling step used to obtain $\mathcal{S}^{\prime}$, and 2) uncertainty in the estimation of the statistic due to finite data. A good estimator should have low variability across repeated sampling of a dataset. In this section we measure both sources of variability and compare the reliability of CCE to ECE.
We perform a large-scale study where we train up to seven different deep regression models on each of the image datasets described in Section 4.1, resulting in 25 distinct trials. In each trial, we measure variability due to MC sampling on a fixed dataset (source 1), by setting $\ell=1$ and repeating the $\overline{CCE}$ computation 5 times to get the mean and standard deviation, $\sigma_{\ell}^{(CCE)}$. To measure the variability of CCE when applied to different evaluation sets (source 2), we approximate its sampling distribution via the bootstrap [13]. We repeatedly resample each dataset (with replacement) to obtain 10 distinct datasets, each of size $n$. For each bootstrapped dataset, we compute $\overline{CCE}$ 5 times (with $\ell=1$ as before). This yields 50 total draws of $\overline{CCE}$ for each dataset-model pair, and we record the standard deviation across draws as $\sigma_{b}^{(CCE)}$. This measurement is a comprehensive summary of the variability of CCE, encompassing both MC sampling and finite data. For comparison, we also compute ECE on each bootstrap to obtain $\sigma_{b}^{(ECE)}$.
Since CCE evaluates a conditional distribution in high dimensions, one might conjecture that sparsity in some input regions could yield higher variance across datasets than what is typically observed in statistics estimated with finite data. However, we find that the variability of CCE, $\sigma_{b}^{(CCE)}$, is statistically indistinguishable from the variability of ECE, $\sigma_{b}^{(ECE)}$. For each dataset-model trial described in the previous paragraph (indexed by $i$), we compute the coefficient of variation, $CV_{i}=\frac{\sigma_{b,i}}{\mu_{b,i}}$, of both CCE and ECE. This value is useful for comparing the variability of data that may live on different scales. Intuitively, a large $CV$ value corresponds to greater variability and therefore less reliability. Letting $\mu_{CV}^{(\cdot)}$ denote the mean CV for a given estimator $(\cdot)$, we first investigate (by way of a two-sample t-test) the hypothesis that CCE has worse reliability than ECE: $H_{0}:\mu_{CV}^{(CCE)}=\mu_{CV}^{(ECE)}$, $H_{a}:\mu_{CV}^{(CCE)}>\mu_{CV}^{(ECE)}$. This test yields a $p$-value of 0.811, causing us to fail to reject the null hypothesis that they are equal. Second, we test the case that the two are different on either side: $H_{a}:\mu_{CV}^{(CCE)}\neq\mu_{CV}^{(ECE)}$. Again, this yields an insignificant $p$-value of 0.378, and we fail to reject the null. The results of our hypothesis tests are summarized in Table 1. Despite its use of Monte Carlo sampling and high-dimensional internals, the reliability of CCE appears to be no different than common calibration estimators used in practice, like ECE.
4.6 Evaluating the robustness of CCE
In this section, we compare the robustness properties of CCE with NLL, another point-wise measurement of probabilistic fit. A known deficiency of NLL is that can be highly sensitive to outliers. In contrast, CCE offers a smoother metric by incorporating similar inputs and outputs to estimate congruence. To demonstrate this behavior, we assess the conditional distributions learned by a Gaussian DNN trained on AAF in two ways: by computing the NLL for each test point, and by computing the CCE for each test point (using the validation split as $\mathcal{S}$). We use t-SNE [43] to project the embeddings of each input $\varphi(x_{i})$ to two dimensions for easy visualization, and create contour plots indicating NLL and CCE across the input space. To build intuition around the geometric structure of the input space, we also provide a plot showing the magnitude of the labels for each test input.
Results are displayed in Figure 4. As expected, we observe that NLL is sensitive to outliers, as evidenced by the concentration of high NLL around a small set of points located near (-40, -18) and (-10, 10). Most other points have roughly equal NLL, with values ranging between 3 and 6. In contrast, CCE is much more robust to these outlier points since it can pool together similar data when estimating congruence. CCE describes a much smoother error surface around local regions of the input space, thus providing clearer insights into the model’s weaknesses. Interestingly, several regions of high CCE correspond to large $y$ values (ages) in the test data. We provide a case study of specific input points in Section 4.7. Overall, we conclude that CCE is more robust to outliers than alternative point-wise calibration measures such as NLL.
0.0161 (0.004)
0.0173 (0.006)
0.0174 (0.007)
0.0179 (0.009)
0.0184 (0.008)
0.1432 (0.025)
0.1282 (0.018)
0.1258 (0.025)
0.1243 (0.019)
0.1220 (0.026)
Figure 5: Test images that incurred the lowest (first row) and highest (second row) CCE values for a $\beta$-Gaussian model [40] trained on the AAF dataset. CCE values indicated above each image. Higher CCE implies a larger discrepancy between the learned and actual conditional distributions.
4.7 Diagnosing point-wise failures without labels
One of the greatest virtues of CCE is that it can estimate point-wise reliability without ground truth labels. Specifically, by viewing the inputs that incur both the highest and lowest CCE values, we can gain insight into a model’s performance on specific regions of $\mathcal{X}$. To do this, we take $\mathcal{S}$ to be the validation split of a dataset (label access), and evaluate $\text{CCE}(f,\mathcal{S},\varphi,\cdot)$ at all inputs from the test split (no label access). We demonstrate this diagnostic process on a $\beta$-Gaussian network, which has been fit on the AAF dataset (see Figure 5). More examples are provided in Appendix E.
In the first row of Figure 5, we plot the images from AAF with the lowest CCE values in our test data. These are images where the model’s learned conditional distributions are most congruent with the true data-generating distribution. In the second row, we plot the images with the highest CCE values, where the model’s distributions are least congruent. The $\beta$-Gaussian network is well-fit on instances of young Asian faces, but struggles when presented with images of older, Caucasian faces. For context, the vast majority of individuals photographed in AAF are of Asian descent. It is possible that CCE has detected a model bias (learned from data) along racial and gender lines.
(a) The effect of the RBF hyperparameter $\gamma$. Red ellipses indicate the region around each point where kernel-implied similarity is meaningfully greater than zero.
(b) The effect of the kernel function class on MCMD. Many choices of kernel function detect similar patterns of regional alignment between the two distributions.
Figure 6: Studying the impact of the kernel functions $k_{\mathcal{X}}$ and $k_{\mathcal{Y}}$ on the MCMD. We vary both kernel hyperparameters (a) and the class of kernel function used (b).
5 Hyperparameter study
In this section, we study the impact of several hyperparameters on different stages of the CCE computation.
First, in Figure 6(a), we show how the estimated MCMD between two distributions (samples generated synthetically) changes when we use an RBF kernel for $k_{\mathcal{X}}$ and $k_{\mathcal{Y}}$ and vary its $\gamma$ parameter. This variation modifies the “radius of influence” for each sample, determining which other input-output pairs contribute to the estimated discrepancy at that point. When $\gamma$ grows large, discrepancies are evaluated at each point in near-isolation, resulting in a bumpier MCMD curve that is more sensitive to individual samples. Meanwhile, when $\gamma$ is small, the radius of influence grows and the MCMD becomes smooth. As mentioned in Section 3.2, we observe consistently reasonable CCE behavior when setting $\gamma=\frac{1}{2\hat{\sigma_{y}^{2}}}$, where $\hat{\sigma_{y}^{2}}=\frac{1}{n-1}\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}$ is estimated from evaluation data.
Next, in Figure 6(b), we investigate the relationship between the MCMD and the class of kernel function. Using the same distributional samples as Figure 6(a), we compute and plot MCMD using three separate kernels: RBF, laplacian, and polynomial. For simplicity, we set $k_{\mathcal{X}}=k_{\mathcal{Y}}$. Despite slightly different MCMD curves, we observe that each kernel function yields similar regional patterns of estimated discrepancy, correctly identifying that the samples are most misaligned near $[1,3]$ and gradually approach distributional equivalence as $x\rightarrow 6$. This result gives us confidence that a specific choice of kernel function will not corrupt our CCE estimates.
Finally, in Figure 7, we demonstrate the effect of $\ell$, which is the number of MC samples we draw from a model’s predictive distributions to form $\mathcal{S}^{\prime}$. For three separate models trained on AAF, and for each value in the grid $\ell\in\{1,...,5\}$, we compute $\overline{\mbox{CCE}}$ three times and record the mean and standard deviation (error bars in the figure indicate $\pm 1$ standard deviation). Though the exact value of $\overline{\mbox{CCE}}$ can vary due to randomness in MC sampling, this variation is small (as first demonstrated in Section 4.5) and does not seem to depend on $\ell$. Moreover, such dispersion does not change the implied ranking of models by probabilistic fit. Since higher values of $\ell$ demand greater computational resources without demonstrably reducing variance, we recommend setting $\ell=1$, as is done throughout this paper.
Additional details and hyperparameter experiments, including an exploration of the regularizer $\lambda$, can be found in Appendix A.5.
Figure 7: The effect of the number of Monte Carlo samples, $\ell$, on $\overline{\mbox{CCE}}$ for three models trained on AAF. For each model we vary the number of samples and compute $\overline{\mbox{CCE}}$ three times. Importantly, we observe that CCE is robust to the number of Monte Carlo samples; this effect is consistent across models.
6 Conclusion
In this paper, we identify issues with current approaches that are commonly used to evaluate the quality of predictive distributions produced by neural networks. Existing measurements like ECE that target calibration are inherently marginal and ignore conditional dependencies in the data, while likelihood-based metrics like NLL suffer from several pitfalls that render them less helpful for point-wise evaluation. We introduce a stronger condition than calibration for assessing probabilistic fit, called congruence, and propose Conditional Congruence Error (CCE) as a novel measure. CCE is a conditional, rather than marginal, statistic and can be computed point-wise for unseen test instances, even in the absence of labels. Through extensive experiments, we show that CCE exhibits several desirable properties including correctness, monotonicity, reliability, and robustness. Future work should study open questions about using CCE in the classification setting, as a loss regularizer, or as an acquisition function in an active learning framework.
Appendix A Additional results
In this section, we provide the results of additional experiments that further support our claims presented in the main body of the paper.
A.1 Full table of results for reliability study
For space considerations, we omit the full experimental results for the variability study in Section 4.5, placing it here (in the supplement) instead. We list results across all trials (dataset, regressor pairs) in Table 2. Note that we do not train Poisson DNN, NB DNN or DDPN on the EVA dataset because the labels in this task are continuous and all three of these models assume discrete likelihoods.
In Table 2, we report MAE to quantify the mean fit of each model on the respective dataset. We also report $\overline{\mbox{CCE}}$, ECE and NLL as measures of probabilistic alignment. The uncertainty of these measurements is quantified with the bootstrap approximation of the sampling distribution as described in Section 4.5, and reported as $\sigma_{l}^{(CCE)}$, $\sigma_{b}^{(CCE)}$, $\sigma_{b}^{(ECE)}$, and $\sigma_{b}^{(NLL)}$. Finally, for $\overline{\mbox{CCE}}$ and ECE we report the coefficient of variation of each trial: $CV_{i}=\frac{\sigma_{b,i}}{\mu_{b,i}}$.
Table 2: Full results of the variability study across all data sets and models.
AAF Model MAE $\overline{\mbox{CCE}}$ $\sigma_{\ell}^{(CCE)}$ $\sigma_{b}^{(CCE)}$ $CV^{(CCE)}$ ECE $\sigma_{b}^{(ECE)}$ $CV^{(CCE)}$ NLL $\sigma_{b}^{(NLL)}$
Poisson DNN 5.325 0.053 0.002 0.004 0.082 0.038 0.007 0.172 3.334 0.025
NB DNN 5.248 0.055 0.002 0.006 0.111 0.034 0.005 0.134 3.340 0.024
Gaussian DNN 5.606 0.052 0.005 0.005 0.069 0.033 0.002 0.047 3.428 0.033
Faithful Gaussian 5.302 0.067 0.002 0.005 0.078 0.099 0.002 0.019 4.606 0.048
Natural Gaussian 6.773 0.077 0.004 0.006 0.079 0.037 0.002 0.060 3.553 0.026
$\beta$-Gaussian 3.254 0.062 0.005 0.004 0.066 0.119 0.006 0.053 3.066 0.039
DDPN 6.207 0.070 0.003 0.006 0.086 0.025 0.006 0.231 3.397 0.015
EVA Gaussian 0.779 0.171 0.011 0.011 0.063 0.179 0.010 0.054 1.526 0.035
Faithful Gaussian 0.738 0.131 0.013 0.009 0.067 0.121 0.009 0.074 1.407 0.040
Natural Gaussian 0.634 0.116 0.008 0.007 0.065 0.121 0.011 0.094 1.163 0.024
$\beta$-Gaussian 0.728 0.113 0.005 0.008 0.072 0.109 0.007 0.067 1.609 0.070
FG-Net Poisson DNN 4.584 0.160 0.013 0.019 0.119 0.053 0.018 0.345 3.077 0.129
NB DNN 3.577 0.150 0.010 0.021 0.141 0.039 0.010 0.252 2.783 0.132
Gaussian DNN 8.600 0.279 0.051 0.036 0.129 0.086 0.013 0.146 3.423 0.049
Faithful Gaussian 4.451 0.188 0.017 0.026 0.135 0.118 0.023 0.196 3.416 0.130
Natural Gaussian 6.152 0.185 0.015 0.027 0.148 0.097 0.021 0.212 3.877 0.250
$\beta$-Gaussian 4.332 0.183 0.018 0.018 0.097 0.094 0.014 0.153 3.849 0.301
DDPN 5.525 0.257 0.017 0.035 0.138 0.234 0.016 0.069 3.761 0.129
COCO-People Poisson DNN 1.121 0.053 0.002 0.002 0.119 0.082 0.003 0.046 1.792 0.013
NB DNN 1.167 0.056 0.002 0.002 0.044 0.084 0.003 0.038 1.767 0.011
Gaussian DNN 1.149 0.030 0.002 0.004 0.043 0.044 0.002 0.032 1.707 0.104
Faithful Gaussian 1.070 0.052 0.001 0.003 0.055 0.070 0.003 0.038 1.928 0.023
Natural Gaussian 1.164 0.040 0.001 0.002 0.046 0.085 0.003 0.031 1.645 0.057
$\beta$-Gaussian 1.057 0.023 0.002 0.003 0.117 0.044 0.003 0.079 1.848 0.071
DDPN 1.135 0.022 0.001 0.002 0.067 0.250 0.002 0.009 1.703 0.106
A.2 Additional monotonicity experiments
In Figure 8 we display the results of our monotonicity experiments on two additional image regression datasets, EVA and FG-Net.